hailees3622:

Convert 11.42424242 … to a rational expression in the form of a over b, where b ≠ 0. i dont know how to convert 11.42424242 to a rational expression

HighIQReptilian:

What they are asking you to do is find the division problem that created this decimal.

hailees3622:

no there asking to convert it into the form a over b

HighIQReptilian:

A over b is a division problem. :)

hailees3622:

oh yh

HighIQReptilian:

The way to do this is to remove the repeating section.

hailees3622:

how do you do that exactly?

HighIQReptilian:

This is done by setting the original number equal to x. In this case, x= 11.4242.... then multiply each side by 100 to get 100x=1142.4242. Then subtract x from each side. (x=11.4242). This will give you 99x=1130.9998. Divide each side by 99 to get x=1130.9998/99. Then simplify this to get the final fraction.

Florisalreadytaken:

i am not following that solution one thing to notice is that in \( 11.4242... \) , the \( .42 \) is periodic, which the whole thing you write as \( 11.\overset{—}{42} \) -- normally, if 42 was a natural number as a fraction you would write it \( \frac{42}{100} \) right? BUT since its periodic, you should always write it over \( n-1 \) so in our case \( 100-1 \) thus, your answer would be \(11\frac{42}{99} \overset{\text{example: }\: 1\frac{2}{3}=\frac{1\cdot \:3+2}{3}}{\ \ \ \ =========\Rightarrow \ \ \ } \frac{1131}{99} \) can we simplify that?

HighIQReptilian:

Your answer seems interesting. It does seem easier to do than the one I provide. However, in the end, they are almost the same with less than a 0.1 difference.

surjithayer:

x=11.424242... x=11+0.424242... x=11+0.42+.0042+.000042+... x=11+y y=0.42+.0042+.000042+... it is a geometric seris a1=0.42 \[r=\frac{ 0.0042 }{ 0.42 }=\frac{ 0.0042 }{ 0.4200 }=\frac{ 1 }{ 100 }<1\] \[y=\frac{ a }{1-r }=\frac{ 0.42 }{ 1-0.01 }=\frac{ 0.42 }{ 0.99}=\frac{ 42 }{ 99 }=\frac{ 14 }{ 33 }\] \[x=11+\frac{ 14 }{ 33 }=\frac{ 363+14 }{ 33 }=\frac{ 377 }{ 33 }\]

surjithayer:

second method let x=11.424242...(1) 10x=114.24242... 100 x=1142.424242... (2) (2)-(1) gives 99x=1131 \[x=\frac{ 1131 }{ 99 }=\frac{ 377 }{ 33 }\]